OpenStudy (anonymous):

How do I solve this and whats the correct answer? 12\[\div\]6\[\div\](-2)(-1)-8+(-6)-(-1)

jimthompson5910 (jim_thompson5910):

\[\Large 12\div 6\div (-2)(-1)-8+(-6)-(-1)\] \[\Large 2\div (-2)(-1)-8+(-6)-(-1)\] \[\Large -1(-1)-8+(-6)-(-1)\] \[\Large 1-8-6+1\] \[\Large -12\] So \[\Large 12\div 6\div (-2)(-1)-8+(-6)-(-1)=-12\]

OpenStudy (anonymous):

12/2/(-1)(-2)-8+(-6)-(-1) =2/(-16)+(-5) =2/(-21) = (-0.0952)